# Multimodal Inverse Attention Network with Intrinsic Discriminant Feature Exploitation for Fake News Detection

### v1

@@ 3.4 Inverse Attention
Exploring intrinsic relationships between modalities remains a significant challenge in multimodal fake news detection, particularly for cases involving mismatched textual and visual content. While the widely adopted cross-attention mechanism has shown promise in modeling cross-modal interactions, it often falls short in effectively capturing inconsistencies between cross-modal features.
Specifically, cross-attention calculates the similarity between query and key vectors to identify which key vectors each query vector should attend to, and then applying the resulting similarity weights to compute a weighted sum of value vectors. While this mechanism effectively emphasizes similar content across modalities, it may inadvertently enable the network to learn implicit pseudo-consistencies, leading to confusion in fake news detection. To address this limitation, we propose the inverse attention mechanism, which amplifies dissimilar features to enhance the learning of explicit inconsistency signals. The detailed explanation of the inverse attention mechanism employed in CIM is as follows:
$\boldsymbol{Att}_{t\to o}^{co-ic}=softmax\left({{\boldsymbol{A}}-{\bf{\boldsymbol{Att}}}_{t\to o}^{co}}\right)$ (14)
$\boldsymbol{\bar{R}}_{t\to o}^{co-ic}=\boldsymbol{Att}_{t\to o}^{co-ic}{\boldsymbol{V}_{t}^{co}}$ (15)
@@
$g=\sigma\left({{\boldsymbol{W}_{g}}\left[{\boldsymbol{\bar{R}}_{t\to o}^{co},\boldsymbol{\bar{R}}_{t\to o}^{co-ic}}\right]+{b_{g}}}\right)$ (16)
$\boldsymbol{R}_{t\to o}^{co}=g\cdot\boldsymbol{\bar{R}}_{t\to o}^{co}+\left({1-g}\right)\cdot\boldsymbol{\bar{R}}_{t\to o}^{co-ic}$ (17)
@@
Given that inconsistencies can also arise within the content of fake news texts or images, the inverse attention mechanism is integrated into both the Local-to-Global and Local-to-Local modules. Specifically, during the computation of $\boldsymbol{R}_{t}^{ll}$, $\boldsymbol{R}_{o}^{ll}$, $\boldsymbol{R}_{t}^{lg}$, and $\boldsymbol{R}_{o}^{lg}$, similar operations to those described in Equations 14 and 15 are applied. This strategy enhances the model’s ability to learn inconsistency features both in intra- and inter-modality, thereby improving its effectiveness in detecting various types of fake news.
Notably, since interactions are performed at the token level, the values of individual elements directly influence the computation of inconsistency weights, particularly when processing textual embeddings. To mitigate the impact of irrelevant or missing information, we introduce a positional mask. Furthermore, due to the potential ambiguity in expression arising from the sequence order of text and the positional information of objects in images, and since the attention mechanism lacks an inherent notion of sequence order, we also incorporate positional encoding information. The specific formulation is as follows:
$P{E_{\left({pos,2i}\right)}}=\sin\left({\frac{{pos}}{{{{10000}^{\frac{{2i}}{d}}}}}}\right)$ (18)
$P{E_{\left({pos,2i+1}\right)}}=\cos\left({\frac{{pos}}{{{{10000}^{\frac{{2i}}{d}}}}}}\right)$ (19)
where, $pos$, $i$ and $d$ represent the position index, dimension index, and the dimension of the positional encoding, respectively.
@@
Table 1: Results of comparison among different approaches on Weibo17, Weibo21, GossipCop and PolitiFact Datasets. The best performance is highlighted in bold and the follow-up is highlighted in underlined.
@@ 4.1 Configurations
Dataset. Our experiments are evaluated on four real-world datasets: Weibo17, Weibo21, GossipCop, and PolitiFact, which cover almost all publicly available datasets used for supervised learning. These datasets used in the experiments not only cover multiple domains but also include specific domain datasets where fake news frequently occurs. The detailed descriptions of these datasets are referred to Appendix A.
Baselines. To evaluate the effectiveness and performance of the proposed method in multimodal fake news detection. The baseline methods include classic models that leverage deep neural networks for extraction such as EANN Wang et al. (2018) and Spotfake Singhal et al. (2019), as well as methods exploring cross-modal relationships such as CAFE Chen et al. (2022), CMC Wei et al. (2022), BMR Ying et al. (2023) and MSACA Wang et al. (2024a), and integrating external knowledge for enhanced detection accuracy such as MRHF Wu et al. (2023). The detailed descriptions of these methodes are referred to Appendix B.
Implementation. All experiments were implemented using the Pytorch toolkit on an NVIDIA A100 GPU. We pad the sequence length of news text to 196 for all datasets, denoted as $m=196$, and put the processed text into the BERT [^1] https://huggingface.co/google-bert/bert-base-uncased. Images were resized to ${224\times 224}$ pixels and subsequently divided into ${u=14\times 14}$ patches, which were then input into the vit-patch16-224222https://huggingface.co/google/vit-base-patch16-224. The multi-head attention used in both the Local-to-Local Block and the Cross-modal Interaction Module consists of 2 layers with 12 heads. The initial learning rate was set to $2{e^{-6}}$, and we utilized the StepLR decay strategy with a decay step of 20 and a decay rate of 0.5 to help the model converge better.
@@ 4.3 Ablation Studies
Impact of each component. In order to address Q2, we evaluate MIAN against its variants by systematically removing specific components. Specifically, w/o intra-lg refers to eliminating the Local-to-Global Block, thereby removing the interaction between global and local features within both text and image branches. w/o intra-lg-ic and w/o intra-ll-ic refer to removing the inconsistency features generated by inverse attention at different levels within the same modality (Local-to-Global and Local-to-Local, respectively). w/o inter-ic refers to excluding inconsistent information between modalities.
Table 2 presents the results of the ablation experiments comparing various method variants, demonstrating the following key findings: (1) In the multimodal fake news detection task, comprehensive and in-depth unimodal information is essential. Extracting features from text or images aids the model in understanding the news content more effectively. (2) The loss of inconsistency information at any stage significantly impacts the accuracy of the detection results, particularly for fake news. Other metrics and results on the GossipCop dataset can be found in the Appendix C.1.
@@ C.2 Experiments on Fakeddit
To further validate the effectiveness of our method, we introduced the Fakeddit dataset. Unlike the common datasets mentioned earlier, where each news article is only labeled as either real or fake, the Fakeddit dataset not only has labels indicating the authenticity of the news but also includes specific labels for five different types of fake news. These labels are True, Satire/Parody, Misleading Content, Imposter Content, and False Connection and Manipulate Content. Here is a brief explanation of each label:
@@
Tables 6 and figure 5 present the comparative and ablation experiment results of our approach across all categories in the Fakeddit dataset. In the comparative experiments, the proposed method achieved the best performance in the categories of True, Satire/Parody, Misleading Content, False Connection, and Manipulated Content, and second-best in the Imposter Content category. This demonstrates that MIAN has strong generalization ability across different types of fake news. In the ablation experiments, other model variants exhibited varying degrees of accuracy degradation, indicating that our proposed fake news classification approach effectively covers various fake news categories, and the introduced modules contribute to improving the model’s performance.
The t-SNE visualizations of the features extracted by MIAN and other methods (CAFE, HMCAN, and LII) on the Fakeddit dataset, as shown in Figure 6, highlight the superior performance of MIAN in separating the six classes of news. Unlike the other methods, MIAN demonstrates clear and distinct clustering, with minimal overlap between real and fake news categories. This indicates that MIAN effectively captures discriminative features across modalities, leading to improved classification accuracy. In contrast, CAFE, HMCAN, and LII exhibit more significant overlap, especially in distinguishing fake news types, suggesting that MIAN’s approach to hierarchical learning and cross-attention is more effective in handling the complexities of multimodal fake news detection.
@@
Table 6: Comparison of accuracy of different models on Fakeddit dataset.

### v2

@@ 3.3 Inverse Attention
Exploring intrinsic relationships between modalities remains a significant challenge in multimodal fake news detection, especially in cases involving mismatched textual and visual content. Although the widely adopted cross-attention mechanism has shown promise in modeling cross-modal interactions, it often fails to capture inconsistencies between cross-modal features effectively.
Specifically, cross-attention calculates the similarity between query and key vectors to determine which key vectors each query vector should attend to, and then applies the resulting similarity weights to compute a weighted sum of value vectors. While this mechanism effectively emphasizes similar content across modalities, it may inadvertently encourage the network to learn implicit pseudo-consistencies, leading to confusion in fake news detection. To address this limitation, we propose the inverse attention mechanism, which amplifies dissimilar features to enhance the learning of explicit inconsistency signals. The detailed explanation of the inverse attention mechanism employed in CIM is as follows:
$\boldsymbol{Att}_{t\to o}^{co-ic}=softmax\left({{\boldsymbol{A}}-{\bf{\boldsymbol{Att}}}_{t\to o}^{co}}\right),$ (14)
$\boldsymbol{\bar{R}}_{t\to o}^{co-ic}=\boldsymbol{Att}_{t\to o}^{co-ic}{\boldsymbol{V}_{t}^{co}},$ (15)
@@
$g=\sigma\left({{\boldsymbol{W}_{g}}\left[{\boldsymbol{\bar{R}}_{t\to o}^{co},\boldsymbol{\bar{R}}_{t\to o}^{co-ic}}\right]+{b_{g}}}\right),$ (16)
$\boldsymbol{R}_{t\to o}^{co}=g\cdot\boldsymbol{\bar{R}}_{t\to o}^{co}+\left({1-g}\right)\cdot\boldsymbol{\bar{R}}_{t\to o}^{co-ic},$ (17)
@@
Given that inconsistencies may also emerge within the content of fake news texts or images, the inverse attention mechanism is incorporated into both the Local-to-Global and Local-to-Local modules. Specifically, during the computation of $\boldsymbol{R}_{t}^{ll}$, $\boldsymbol{R}_{o}^{ll}$, $\boldsymbol{R}_{t}^{lg}$, and $\boldsymbol{R}_{o}^{lg}$, similar operations to those described in Equations 14 and 15 are applied. This design enables the model to capture inconsistent features at both intra- and inter-modal levels, thereby improving its effectiveness in detecting various types of fake news.
Notably, since interactions are performed at the token level, the values of individual elements directly influence the computation of inconsistency weights, particularly when processing textual embeddings. To alleviate the impact of irrelevant or missing information, we introduce a positional mask. Furthermore, due to the potential ambiguity in expression arising from the sequence order of text and the positional information of objects in images, and since the attention mechanism lacks an inherent notion of sequence order, we also incorporate positional encoding. The specific formulation is as follows:
$P{E_{\left({pos,2i}\right)}}=\sin\left({\frac{{pos}}{{{{10000}^{\frac{{2i}}{d}}}}}}\right),$ (18)
$P{E_{\left({pos,2i+1}\right)}}=\cos\left({\frac{{pos}}{{{{10000}^{\frac{{2i}}{d}}}}}}\right),$ (19)
where $pos$, $i$ and $d$ represent the position index, dimension index, and the dimension of the positional encoding, respectively.
@@
Table 1: Results of comparison among different approaches on Weibo17, Weibo21, GossipCop and PolitiFact Datasets. The best performance is highlighted in bold, and the follow-up is highlighted in underlined.
@@ 4.1 Configurations
Dataset. Our experiments are evaluated on four real-world datasets: Weibo17, Weibo21, GossipCop, and PolitiFact, which cover almost publicly available datasets. These datasets used in the experiments not only cover multiple domains but also include specific domain datasets where fake news frequently occurs. The detailed descriptions of these datasets are referred to Appendix A.
Baselines. To evaluate the effectiveness and performance of the proposed method in multimodal fake news detection. The baseline methods include classic models that leverage deep neural networks for extraction such as EANN Wang et al. (2018), Spotfake Singhal et al. (2019), HMCAN Qian et al. (2021) and RaCMC Yu et al. (2025), as well as methods exploring cross-modal relationships such as CAFE Chen et al. (2022), CMC Wei et al. (2022), BMR Ying et al. (2023) and MSACA Wang et al. (2024a), and integrating external knowledge for enhanced detection accuracy such as MRHF Wu et al. (2023). The detailed descriptions of these methods are referred to Appendix B.
Implementation. All experiments were implemented using the PyTorch toolkit on an NVIDIA A100 GPU. We pad the sequence length of news text to 196 for all datasets, denoted as $m=196$, and put the processed text into the BERT. Images were resized to ${224\times 224}$ pixels and subsequently divided into ${u=14\times 14}$ patches, which were then input into the vit-patch16-224. The multi-head attention used in both the Local-to-Local Block and the Cross-modal Interaction Module consists of 2 layers with 12 heads. The initial learning rate was set to $2{e^{-6}}$, and we utilized the StepLR decay strategy with a decay step of 20 and a decay rate of 0.5 to help the model converge better.
@@ 4.3 Ablation Studies
Impact of each component. In order to address Q2, we evaluate MIAN against its variants by systematically removing specific components. Specifically, w/o intra-lg refers to eliminating the Local-to-Global Block, thereby removing the interaction between global and local features within both text and image branches. w/o intra-lg-ic and w/o intra-ll-ic refer to removing the inconsistent features generated by inverse attention at different levels within the same modality (Local-to-Global and Local-to-Local, respectively). w/o inter-ic refers to excluding the inconsistent features in the CIM.
Table 2 presents the results of the ablation experiments comparing various method variants, demonstrating the following key findings: (1) In the multimodal fake news detection task, comprehensive and in-depth unimodal information is essential. Extracting features from text or images aids the model in understanding the news content more effectively. (2) The loss of inconsistent information at any stage significantly impacts the accuracy of the detection results, particularly for fake news. Other metrics and results on the GossipCop dataset can be found in the Appendix C.1.
@@ 5 Conclusion & Limitations
Acknowledgments
This work was supported by the Joint Project for Innovation and Development of Shandong Natural Science Foundation, (No. ZR2022LZH012).
Appendix
@@ C.1 The detailed experiments
To further validate the reliability of our results, we conducted statistical significance tests to assess whether the observed performance improvements of MIAN over baseline methods are consistent and non-random. As shown in Table 5, the results confirm that the improvements are statistically significant, supporting the robustness of our conclusions and reinforcing the credibility of the comparative evaluations.
Dataset p-value (acc) p-value (F-F1) p-value (R-F1)
Weibo17 0.0254* 0.0184* 0.0151*
Weibo21 0.0453* 0.0619 0.0279*
GossipCop 0.0002*** 0.0001*** 0.0006***
PolitiFact 0.0042** 0.0377* 0.0117*
Table 5: Significance level is marked with *: significant (p-value $<$ 0.05) , **: very significant (p-value$<$0.01), ***: highly significant (p-value$<$0.001). Results without a mark are not significant.
@@ C.2 Experiments on Fakeddit
To further validate the effectiveness of our method, we introduced the Fakeddit dataset. Unlike the common datasets mentioned earlier, where each news article is only labeled as either real or fake, the Fakeddit dataset not only has labels indicating the authenticity of the news but also includes specific labels for five different types of fake news. These labels are True, Satire/Parody, Misleading Content, Imposter Content, False Connection, and Manipulate Content. Here is a brief explanation of each label:
@@
Table 7 and Figure 5 present the comparative and ablation experiment results of our approach across all categories in the Fakeddit dataset. In the comparative experiments, the proposed method achieved the best performance in the categories of True, Satire/Parody, Misleading Content, False Connection, and Manipulated Content, and second-best in the Imposter Content category. This demonstrates that MIAN has strong generalization ability across different types of fake news. In the ablation experiments, other model variants exhibited varying degrees of accuracy degradation, indicating that our proposed fake news classification approach effectively covers various fake news categories, and the introduced modules contribute to improving the model’s performance.
The t-SNE visualizations of the features extracted by MIAN and other methods (CAFE, HMCAN, and LII) on the Fakeddit dataset, as shown in Figure 6, highlight the superior performance of MIAN in separating the six classes of news. Unlike the other methods, MIAN demonstrates clear and distinct clustering, with minimal overlap between real and fake news categories. This indicates that MIAN effectively captures discriminative features across modalities, leading to improved classification accuracy. In contrast, CAFE, HMCAN, and LII exhibit more significant overlap, especially in distinguishing fake news types, suggesting that MIAN’s approach to hierarchical learning and co-attention is more effective in handling the complexities of multimodal fake news detection.
@@
Table 7: Comparison of the accuracy of different models on Fakeddit dataset.